Please help me solve this equation! log2 x+ log2 (x+1)=1
Answer by anshu_a(16)

log2 x + log2 (x+1) = 1
log2 (x*(x+1)) = 1
x(x+1) = 2^1 , x^2-x-2 = 0
(x+2)(x-1) = 0
x = 1,-2